OpenStudy (anonymous):

Are these planes parallel and distinct or parallel and coincident ? 3x+9y-6z-24=0 4x+6y-14z-8=0

OpenStudy (tkhunny):

Divide the first by 3 and the second by 2, for starters.

OpenStudy (anonymous):

@tkhunny and then ??

OpenStudy (tkhunny):

What do you get?

OpenStudy (anonymous):

x=4+5z

OpenStudy (tkhunny):

?? What is that? I was hoping you would get: x+3y-2z-8=0 2x+3y-7z-4=0

OpenStudy (anonymous):

Yes i did

OpenStudy (anonymous):

Then i did elimination to get rid of y

OpenStudy (tkhunny):

Why? Are they parallel?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

We have to figure of if theyre parallel and distinct or coincidentt

OpenStudy (tkhunny):

They are? What makes you think so? When you eliminated y, didn't that give you a line in the x-z plane?

OpenStudy (anonymous):

its given in the question

OpenStudy (tkhunny):

Are you going to buy it, just because it says so? If they are parallel, shouldn't they have the same Normal Vector?

OpenStudy (anonymous):

so they're not parallel, they intersect ?

OpenStudy (tkhunny):

That's what they look like to me.

OpenStudy (anonymous):

Answer key says parallel and coincident

OpenStudy (tkhunny):

3x+9y-6z-24=0 4x+6y-14z-8=0 Find a point on one of them. (10,0,1) Prove 3(10) + 9(0) - 6(1) - 24 = 30 + 0 - 6 - 24 = 24 - 24 = 0 Is it on the other one? 4(10) + 6(0) - 14(1) - 8 = 40 - 14 - 8 = 26 - 8 = 18 \(\ne 0\) Okay, we just proved that they are NOT coincident. Did you write the problem correctly? If you did, we're done.